OpenStudy (anonymous):

7x -4y=20 for y when x=5 ????????????

jigglypuff314 (jigglypuff314):

plug in 5 for x and then simplify and isolate y so 7(5) - 4y = 20 35 - 4y = 20 subtract 35 from both sides -35 -35 - 4y = -15 then divide both sides by -4 y = (-15)/(-4) = 15/4